OpenStudy (anonymous):

The depth of snow after n hours of a snowstorm is represented by the function f(n + 1) = f(n) + 0.8 where f(0) = 2.5. Which statement describes the sequence of numbers generated by the function? The depth of snow was 0.8 inches when the storm began, and 2.5 inches after the first hour of the storm. The depth of snow was 1.7 inches when the storm began, and 0.8 inches of snow fell each hour. The depth of snow was 2.5 inches when the storm began, and increased by 0.8 inches each hour. The depth of snow was 3.3 inches when the storm began, and 2.5 inches of snow fell in 1 hour.

OpenStudy (anonymous):

@chlobohoe

OpenStudy (anonymous):

f(0) gives you the amount when it began and the amount you're adding to f(n) gives the additional snowfall

OpenStudy (anonymous):

so whats the problem i need to solve?

OpenStudy (anonymous):

What number is the initial amount of snow?

OpenStudy (anonymous):

0.8

OpenStudy (anonymous):

Nearly always, time starts at 0, and f(0) = 2.5 so the initial snow is 2.5

OpenStudy (anonymous):

f(n + 1) = f(n) + 0.8 is saying that every hour there is 0.8 inches more snow. So after the first hour it's f(1) = f(0) + 0.8 = 2.5 + 0.8 = 3.3 After the 2nd hour f(2) = f(1) + 0.8 = 3.3 + 0.8 = 4.1 and so on